OpenStudy (anonymous):

Use the technique of completing the square to transform the quadratic equation into the form (x + c)2 = a. 2x2 + 24x + 16 = 0

OpenStudy (anonymous):

what happened with the last one?

OpenStudy (anonymous):

did you finish it?

OpenStudy (anonymous):

yes. and got the correct answer. thanks for the help.

OpenStudy (anonymous):

2x^2 + 24x + 16 = 0

OpenStudy (anonymous):

same idea here \[x^2+12x+8=0\\ x^2+12x=-8\\ (x+6)^2=-8+36\\ 111\]

OpenStudy (anonymous):

forget the \(111\) that was a typo

OpenStudy (anonymous):

lol i was like what the? haha ok.

OpenStudy (anonymous):

(x+6)^2=28???? thats what i got.